# Molecular optomechanics with atomic antennas

### v1

@@ Appendix A Derivation of the interaction Hamiltonian
$\hat{H}_{I,\text{lin}}$ $=-\frac{1}{2}R_{ij}Q^{0}\omega_{\sigma}^{2}\mu_{0}$
×$\times\left[\hat{a}^{\dagger}\alpha_{\sigma}(\mathbf{e}_{a})_{j}(\mathbf{G}(\mathbf{r}_{\sigma},\mathbf{r}_{m},\omega_{\sigma})\cdot\mathbf{d}_{\sigma})^{*}_{i}+\text{h.c.}\right]\left(\hat{b}+\hat{b}^{\dagger}\right).$ (23)
@@
[…]
×$~{}~{}~{}\times\left(\hat{b}+\hat{b}^{\dagger}\right).$ (24)
We can choose the phase of the incident laser, or $\alpha_{\sigma}$, in such a way that $\alpha_{\sigma}\mathbf{e}_{a}\cdot(\mathbf{G}\mathbf{d}_{\sigma})^{*}$ becomes a real number, arriving at the interaction Hamiltonian given in Eq. (I.3).
@@ Appendix B Power, flux and efficiency of Raman scattering
In the main text, we introduced the expressions for the emission spectra as given by the Fourier transform of a two-time correlator, multipled by $\omega^{4}$. This allowed us to trace the general trends of the emission intensities. However, to learn about the efficiencies of the Raman processes, here we carry out a more careful analysis.
@@
$P=$ $2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}$
×$\times\int\text{d}\tau e^{i\omega\tau}\langle\hat{\sigma}^{\dagger}(0)\hat{\sigma}(\tau)\rangle$ (25)
@@
$P=$ $2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}$
×$\times\int\text{d}\tau e^{i\omega\tau}\langle\hat{a}^{\dagger}(0)\hat{a}(\tau)\rangle$ (27)
@@
We now integrate the frequencies around the Stokes emission frequency $\omega_{S}$, using Eq. (I.4) to turn the last part turns of the expression for the power to
@@
$P_{\sigma}=$ $4\varepsilon_{0}c|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin(\theta)r^{2}f(r,\theta,\omega_{S})\int\text{d}\phi$
$=$ $\frac{\omega_{S}^{4}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (30)
@@
$\eta_{\sigma}$ $=\frac{P_{\sigma}/(\hbar\omega_{S})}{I\sigma_{\text{spot}}/(\hbar\omega_{l})}$
$=\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{|g_{\sigma}|^{2}}{I\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$~{}~{}~{}~{}\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (31)
@@
$\eta^{\prime}=$ $\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{1}{\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (33)
@@ Appendix C Derivation of the emission spectrum
Notably, the last step requires that the dynamics described by the integral kernel be much faster than the dynamics of $\hat{\tilde{b}}(s)$ (determined by the effective dissipation rate of the mechanical mode $\Gamma$), allowing us to approximate it as $\hat{\tilde{b}}(t)$. This Markov approximation therefore requires that $\Gamma\ll\kappa$. Another way to think about this approximation is by comparing the width of the emission Stokes/anti-Stokes lines to the spectrum of the cavity into which the molecule emits. For $\Gamma/2\pi=1$ THz, and $\kappa/2\pi=50$ THz, and the condition holds trivially.
@@
$=2\omega^{4}\left[\frac{\eta\omega^{4}|g_{\sigma}|^{2}n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{-i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)+\frac{|g_{\sigma}|^{2}(n_{b}+1)}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)\right]$
@@
$=2\omega^{4}|g_{\sigma}|^{2}\left[\frac{n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}+\omega_{l}-\omega)^{2}+(\Gamma/2)^{2}}+\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}-\omega_{l}+\omega)^{2}+(\Gamma/2)^{2}}\right].$ (42)
@@ Appendix D Notes on the estimates for Table 1
$G_{\text{nf}}(\mathbf{r}_{\sigma},\mathbf{r}_{m})\approx\frac{1}{4\pi k^{2}|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|^{3}}.$ (45)
Finally, as an example of a Raman-active molecule, we consider a $\omega_{b}/2\pi=40.5$ THz ($1351~{}\text{cm}^{-1}$) vibrational mode of the R6G [11] with Raman activity in the Gaussian units given as $3.5\times 10^{2}\varepsilon_{0}^{2}\text{\AA}^{4}\text{amu}^{-1}$, and here assumed to be equal to $R^{2}$; accounting for the transformation between Gaussian and SI units (see the SM of [11]) we find
@@ Appendix E Parameters of conventional plasmonic SERS
    where $Q=\omega_{a}/\kappa$ denotes the quality factor of the mode; this expression differs by a factor of $\sqrt{2}$ from the formulation given in [29]. When decoupled from (or weakly coupled to) the Raman-active molecule, and driven on resonance, the cavity settles into a coherent state with amplitude
@@
Appendix F Details of the numerical simulations
To calculate the characteristics of an atomic antenna-enhanced Raman setup, we have implemented two COMSOL models which share the same geometry: a gold nanoparticle with radius of 100 nm, and permittivity $\varepsilon_{\text{Au}}=-9.53+1.51i$ ([52] at 602 nm), separated by distance $d$ from the flat diamond surface with permittivity $\varepsilon_{\text{diam.}}=n_{\text{diam.}}^{2}=5.86$. The computational domain was a sphere with radius of 3 $\mu$m, with a perfectly matching layer (PML) of 1 $\mu$m thickness. Meshing was chosen to ensure convergence, and increasingly fine towards the gap, where the element size was maxed as 0.2 nm.

### v2

@@ A.2 Derivation of the linearized interaction Hamiltonian
$\hat{H}_{I,\text{lin}}$ $=-\frac{1}{2}R_{ij}Q^{0}\omega_{\sigma}^{2}\mu_{0}\left[\hat{a}^{\dagger}\alpha_{\sigma}(\mathbf{e}_{a})_{j}(\mathbf{G}(\mathbf{r}_{\sigma},\mathbf{r}_{m},\omega_{\sigma})\cdot\mathbf{d}_{\sigma})^{*}_{i}+\text{h.c.}\right]\left(\hat{b}+\hat{b}^{\dagger}\right).$ (30)
@@
$\hat{H}_{I,\text{lin}}$ $=-\frac{1}{2}RQ^{0}\omega_{\sigma}^{2}\mu_{0}\left[\hat{a}^{\dagger}\alpha_{\sigma}\mathbf{e}_{a}\cdot(\mathbf{G}(\mathbf{r}_{\sigma},\mathbf{r}_{m},\omega_{\sigma})\cdot\mathbf{d}_{\sigma})^{*}+\text{h.c.}\right]\left(\hat{b}+\hat{b}^{\dagger}\right).$ (31)
We can choose the phase of the incident laser, or $\alpha_{\sigma}$, in such a way that $\alpha_{\sigma}\mathbf{e}_{a}\cdot(\mathbf{G}\mathbf{d}_{\sigma})^{*}$ becomes a real number, arriving at the interaction Hamiltonian given in Eq. (II.1.3).
@@ Appendix B Power, flux and efficiency of Raman scattering
In the main text, we introduced the expressions for the emission spectra as given by the Fourier transform of a two-time correlator, multiplied by $\omega^{4}$. This allowed us to trace the general trends of the emission intensities. However, to learn about the efficiencies of the Raman processes, here we carry out a more careful analysis.
@@
$P=2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}\int\text{d}\tau e^{i\omega\tau}\langle\hat{\sigma}^{\dagger}(0)\hat{\sigma}(\tau)\rangle$ (32)
@@
$P=2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}\int\text{d}\tau e^{i\omega\tau}\langle\hat{a}^{\dagger}(0)\hat{a}(\tau)\rangle$ (34)
@@
We now integrate the frequencies around the Stokes emission frequency $\omega_{S}$, using Eq. (II.1.4), to turn the last part of the expression for the power to
@@
$P_{\sigma}$ $=4\varepsilon_{0}c|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\int\text{d}\theta\sin(\theta)r^{2}f(r,\theta,\omega_{S})\int\text{d}\phi$
$=\frac{\omega_{S}^{4}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (37)
@@
$\eta_{\sigma}$ $=\frac{P_{\sigma}/(\hbar\omega_{S})}{I\sigma_{\text{spot}}/(\hbar\omega_{l})}=\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{|g_{\sigma}|^{2}}{I\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (38)
@@
$\eta^{\prime}=$ $\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{1}{\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (40)
@@ Appendix C Derivation of the emission spectrum
In this section we derive the emission spectra from the systems under weak, and strong illumination.
C.1 Raman emission in the weakly pumped, linearized regime
@@
Notably, the last step requires that the dynamics described by the integral kernel be much faster than the dynamics of $\hat{\tilde{b}}(s)$ (determined by the effective dissipation rate of the mechanical mode $\Gamma$), allowing us to approximate it as $\hat{\tilde{b}}(t)$. This Markov approximation therefore requires that $\Gamma\ll\kappa$. Another way to think about this approximation is by comparing the width of the emission Stokes/anti-Stokes lines to the spectrum of the cavity into which the molecule emits. For $\Gamma/2\pi=1$ THz, and $\kappa/2\pi=50$ THz, this condition holds trivially.
@@
$=2\omega^{4}\left[\frac{\eta\omega^{4}|g_{\sigma}|^{2}n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{-i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)\right.$
@@
$\quad\quad\quad\quad\left.+\frac{|g_{\sigma}|^{2}(n_{b}+1)}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)\right]$
$=2\omega^{4}|g_{\sigma}|^{2}\left[\frac{n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}+\omega_{l}-\omega)^{2}+(\Gamma/2)^{2}}\right.$
$\quad\quad\quad\quad\quad\quad\left.+\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}-\omega_{l}+\omega)^{2}+(\Gamma/2)^{2}}\right].$ (49)
C.2 Raman emission beyond the linearization regime
In Section II.3, we discuss the emission spectra of the system under strong illumination — in the Mollow regime — where the coherence of the atomic antenna $\left\langle\hat{\sigma}\right\rangle$ rapidly decays with the increasing illumination intensity. Simultaneously, we observe in Fig. 3(c,d), that the intensity of Raman emission appears to saturate. This suggests that the Raman emission becomes decoupled from the atomic coherence, in a clear departure from the linearization picture discussed above, and in Section II.1.3.
To understand this behavior, we note the similarity of our system to that discussed by Neuman and co-authors[51], to treat Surface-Enhanced Resonant Raman Scattering under strong illumination. Adopting the approach used in that work, we start with the interaction Hamiltonian (Eq. (II.1.1)), split the operator $\hat{\sigma}\rightarrow\alpha_{\sigma}+\delta\hat{\sigma}$, and perform the rotating wave approximation to remove terms oscillating at twice the optical frequencies:
[…]
In the absence of any correlations in the optical input noise $\hat{a}_{\text{in}}$, we can neglect that term in writing down the Heisenberg equation for the optical cavity operator $\hat{a}$:
$\frac{\text{d}}{\text{d}t}\hat{a}(t)=\left(-i\Delta_{a}-\frac{\kappa}{2}\right)\hat{a}(t)+ig_{0,\sigma}(\alpha_{\sigma}+\delta\hat{\sigma})(\hat{b}+\hat{b}^{\dagger}),$ (51)
with formal solution
$\hat{a}(t)=$ $~{}ig_{0,\sigma}\int_{-\infty}^{t}e^{(-i\Delta_{a}-\kappa/2)(t-s)}\alpha_{\sigma}(s)[\hat{b}(s)+\hat{b}^{\dagger}(s)]$
$+ig_{0,\sigma}\int_{-\infty}^{t}e^{(-i\Delta_{a}-\kappa/2)(t-s)}\delta\hat{\sigma}(s)[\hat{b}(s)+\hat{b}^{\dagger}(s)].$ (52)
In the Mollow regime, the first term can be made arbitrarily small by choosing sufficiently large laser intensity, and so we focus on the second line of the above expression. Using again operators $\hat{\tilde{b}}(s)$, which evolve at the rate far smaller than $\kappa$, we can perform the Markov approximation and find
$\hat{a}(t)\approx$ $~{}ig_{0,\sigma}\left[\frac{\delta\hat{\sigma}(t)\hat{b}(t)}{-i\Delta_{a}+i\omega_{b}-\kappa/2}+\frac{\delta\hat{\sigma}(t)\hat{b}^{\dagger}(t)}{-i\Delta_{a}-i\omega_{b}-\kappa/2}\right].$ (53)
As before, the emission spectrum from these fluctuations can be calculated from the Fourier transform of the two-time correlator
$\left\langle\hat{a}^{\dagger}(0)\hat{a}(t)\right\rangle\approx$ $\left|\frac{g_{0,\sigma}}{\kappa/2+i(-\Delta_{a}+\omega_{b})}\right|^{2}\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}^{\dagger}(0)\delta\hat{\sigma}(t)\hat{b}(t)\right\rangle$
$+\left|\frac{g_{0,\sigma}}{\kappa/2-i(-\Delta_{a}+\omega_{b})}\right|^{2}\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}(0)\delta\hat{\sigma}(t)\hat{b}^{\dagger}(t)\right\rangle$
$=$ $\frac{g_{0,\sigma}^{2}}{\kappa^{2}/4+(-\Delta_{a}+\omega_{b})^{2}}\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}^{\dagger}(0)\delta\hat{\sigma}(t)\hat{b}(t)\right\rangle$
$+\frac{g_{0,\sigma}^{2}}{\kappa^{2}/4+(-\Delta_{a}-\omega_{b})^{2}}\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}(0)\delta\hat{\sigma}(t)\hat{b}^{\dagger}(t)\right\rangle,$ (54)
where we dropped terms proportional to the squeezing of the mechanical mode. If we additionally factorize the correlators, evaluate the correlators assuming decoupled dynamics, and use correlators given in Eqs. (47) and (48), we find
$\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}^{\dagger}(0)\delta\hat{\sigma}(t)\hat{b}(t)\right\rangle\approx\left\langle\delta\hat{\sigma}^{\dagger}(0)\delta\hat{\sigma}(t)\right\rangle\left\langle\hat{b}^{\dagger}(0)\hat{b}(t)\right\rangle\approx\left\langle\delta\hat{\sigma}^{\dagger}(0)\delta\hat{\sigma}(0)\right\rangle n_{b}^{\mathrm{th}}e^{(-i\omega_{b}-\Gamma/2)t}$
$\left\langle\delta\hat{\sigma}^{\dagger}(0)\hat{b}(0)\delta\hat{\sigma}(t)\hat{b}^{\dagger}(t)\right\rangle\approx\left\langle\delta\hat{\sigma}^{\dagger}(0)\delta\hat{\sigma}(t)\right\rangle\left\langle\hat{b}(0)\hat{b}^{\dagger}(t)\right\rangle\approx\left\langle\delta\hat{\sigma}^{\dagger}(0)\delta\hat{\sigma}(0)\right\rangle(n_{b}^{\mathrm{th}}+1)e^{(i\omega_{b}-\Gamma/2)t}.$ (55)
Finally, the emission spectrum takes on a form very similar to that obtained through the linearization procedure (Eq. (C.1)):
$S_{a}(\omega)$ $=2\omega^{4}g_{0,\sigma}^{2}\left\langle\delta\hat{\sigma}^{\dagger}\delta\hat{\sigma}\right\rangle\left[\frac{n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}+\omega_{l}-\omega)^{2}+(\Gamma/2)^{2}}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}-\omega_{l}+\omega)^{2}+(\Gamma/2)^{2}}\right],$ (56)
except for the change of $|g_{\sigma}|^{2}=g_{0,\sigma}^{2}|\alpha_{\sigma}|^{2}\rightarrow g_{0,\sigma}^{2}\left\langle\delta\hat{\sigma}^{\dagger}\delta\hat{\sigma}\right\rangle$. This observation would suggest that the overall intensity of the Raman scattering should depend explicitly only on the population of the atomic antenna, irrespective of whether it is in a coherent, or a maximally mixed state. We should note that the derivation carried out in this, and the previous subsection, rely on the same approximation, of decoupling the state of the atomic antenna from other components of the system. This approximation would likely break down only for the stronger single-photon coupling rates $g_{0,\sigma}$.
@@ Appendix D Notes on the estimates for Table 1
$G_{\text{nf}}(\mathbf{r}_{\sigma},\mathbf{r}_{m})\approx\frac{1}{4\pi k^{2}|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|^{3}},$ (59)
and use it to estimate the electric field from the antenna at the position of the molecule
$|\mathbf{e}_{\sigma}(\mathbf{r}_{m})|\approx\omega_{\sigma}^{2}\mu_{0}G(\mathbf{r}_{\sigma},\mathbf{r}_{m})|\mathbf{d}_{\sigma}|.$ (60)
Finally, as an example of a Raman-active molecule, we consider a $\omega_{b}/2\pi=40.5$ THz ($1351~{}\text{cm}^{-1}$) vibrational mode of the R6G [11] with Raman activity in the Gaussian units given as $3.5\times 10^{2}\varepsilon_{0}^{2}\text{\AA}^{4}\text{amu}^{-1}$, and here assumed to be equal to $R^{2}$; accounting for the transformation between Gaussian and SI units (see the SM of Ref. [11]) we find
@@ Appendix E Parameters of conventional plasmonic SERS
    where $Q=\omega_{a}/\kappa$ denotes the quality factor of the mode; this expression differs by a factor of $\sqrt{2}$ from the formulation given in Ref. [29]. When decoupled from (or weakly coupled to) the Raman-active molecule, and driven on resonance, the cavity settles into a coherent state with amplitude
@@
Appendix F Numerical modelling
F.1 Modelling the classical response of the setup
To calculate the characteristics of an atomic antenna-enhanced Raman setup, we have implemented two COMSOL models which share the same geometry: a gold nanoparticle with radius of 100 nm, and permittivity $\varepsilon_{\text{Au}}=-9.53+1.51i$ (value taken from Ref. 59, at 602 nm), separated by distance $d$ from the flat diamond surface with permittivity $\varepsilon_{\text{diam.}}=n_{\text{diam.}}^{2}=5.86$. The computational domain was a sphere with radius of 3 $\mu$m, with a perfectly matching layer (PML) of 1 $\mu$m thickness. Meshing was chosen to ensure convergence, and increasingly fine towards the gap, where the element size was maxed as 0.2 nm.
@@ F.1 Modelling the classical response of the setup
In Fig. 5 we plot the Purcell enhancement (solid orange line), split between the absorption (dashed green line) and radiation (dashed blue line), as a function of spacing between the substrate and the nanoparticle.
Figure 5: Numerically calculated Purcell factor (solid orange line), as a function of spacing between the substrate and the nanoparticle (see the schematic in Fig. 2(a) and Section F.1 for details of this model). Absorptive and radiative channels are depicted with the dashed green blue lines, respectively
F.1.1 Derivation of Eq. (19)
In bulk diamond, the total decay rate of the GeV $\gamma=\gamma_{\text{nr}}+\gamma_{\text{r}}$ is determined by its radiation (with rate $\gamma_{\text{r}}=\gamma_{0}$), and nonradiative internal processes (with rate estimated from experiments as $\gamma_{\text{nr}}=3\gamma_{0}$ at 50 K) [26]. When the defect is coupled to a structured environment, the internal nonradiative processes still occur at the same rate $\gamma_{\text{nr}}$. Conversely, the rate of emission from the emitter $\gamma_{\text{r}}$ is enhanced via the Purcell effect to $F_{P}\gamma_{\text{r}}$ (whether this enhancement is related to the increase of the radiative density of states, or absorption in the environment), and the total decay rate becomes
$\gamma\rightarrow\gamma_{\text{tot}}$ $=F_{P}\gamma_{0}\,+3\gamma_{0},$ (69)
which is the first line in Eq. (II.2.1).
In our work, the Purcell factor is numerically calculated as a ratio of the power absorbed in the environment and radiated from an electric dipole (positioned at the atomic antenna, and characterised with the quantum efficiency of 1) with, and without the nanoparticle:
$F_{P}=\frac{(P_{\text{rad}}+P_{\text{abs}})~{}\text{(with nanoparticle)}}{P_{\text{rad}}^{0}~{}\text{(without nanoparticle)}},$ (70)
where in the absence of nanoparticle we explicitly denote radiated power as $P_{\text{rad}}^{0}=P_{0}$, and note that there are no losses in the medium that could yield absorption.
While the Purcell factor could alternatively be calculated from the Green’s function of the system, we found that the former approach is more numerically stable.
F.2 Modelling the quantum dynamics of the system
To understand the response of the system, comprising atomic antenna ($\hat{\sigma}$), plasmonic cavity mode ($\hat{a}$) and Raman-active molecule with a single vibrational mode ($\hat{b}$), in more detail (especially beyond the linearization regime), we numerically model its dynamics using QuTiP solvers [49, 50]. The coherent dynamics is governed by the full Hamiltonian, accounting for all the terms discussed in Section A.1 and in Fig. 4:
$\hat{H}=~{}$ $\hbar\omega_{b}\hat{b}^{\dagger}\hat{b}+\hbar(\omega_{\sigma}-\omega_{l})\hat{\sigma}^{\dagger}\hat{\sigma}+\hbar(\omega_{a}-\omega_{l})\hat{a}^{\dagger}\hat{a}$
$+\hbar\Omega\left(\hat{\sigma}+\hat{\sigma}^{\dagger}\right)+\hbar\Omega_{a}\left(\hat{a}+\hat{a}^{\dagger}\right)$
$-\hbar g_{0,\sigma}\left(\hat{\sigma}^{\dagger}\hat{a}+\hat{a}^{\dagger}\hat{\sigma}\right)\left(\hat{b}+\hat{b}^{\dagger}\right)-\hbar g_{0,a}\hat{a}^{\dagger}\hat{a}\left(\hat{b}+\hat{b}^{\dagger}\right)$
$+\hbar g_{JC}\left(\hat{\sigma}^{\dagger}\hat{a}+\hat{a}^{\dagger}\hat{\sigma}\right).$ (71)
Besides the coherent driving ($\propto\Omega$), and optomechanical coupling mediated by the atomic antenna ($\propto g_{0,\sigma}$), we also include terms describing pumping of the cavity mode ($\propto\Omega_{a}$), conventional Raman interaction ($\propto g_{0,a}$), and the Jaynes-Cummings coupling between the GeV and cavity mode (see discussion in Section A.1).
The incoherent dynamics is described as in the main text (Section II.1.2).
For the calculations discussed in the main text, we use coupling, frequency and dissipation parameters as defined in Table 1 and take $g_{JC}/2\pi=10^{11}$ Hz (which would yield the Purcell factor around 2, identified for the spacings between the substrate and nanoparticle larger than 12 nm; see Section A.1). To calculate the spectra in Fig. 3(a,c), we use the following laser intensities:
[…]
$6.9\times 10^{-2},~{}0.17~{}\upmu\text{W}/\upmu\text{m}^{2};$
[…]
$0.17,~{}1.2,~{}6.9,~{}43.4~{}\upmu\text{W}/\upmu\text{m}^{2}.$
These values are marked with vertical dashed lines in Fig. 3(b,d).
To identify the Raman scattering from conventional SERS and calculate the emission spectra $S_{a,\text{conv}}$, we run the same simulations setting $g_{0,\sigma}=0$.